Jim's age is 4 times that of his sister. The product of their ages is 36. How old is each?
Answer by farah(18)

let J be Jim and S be sister..
J=4S
JS=36
4S(S)=36
4S^2=36
S^2=9
square root of 9 is 3
so, sister's age is 3, and jim's age = 4(3) is 12
sister= 3 years old
jim = 12 years old